OpenStudy (zasharra):

Andrea is designing the seating arrangement for a concert in her local park. To give everyone a good view, each row must have 4 more seats than the row before it, and the first row can only have 10 seats. Help Andrea plan the rest of the seating by solving for how many seats are in row 20. Then explain to Andrea how to create an equation to predict the number of seats in any row. Show your work, and use complete sentences.

OpenStudy (zasharra):

@Directrix

OpenStudy (zasharra):

@robtobey @math&ing001

OpenStudy (isaidavila):

Is like a sequence of numbers, suppose the first row has 1 seat so the next will have 5 and the third 9 and so on 1,5,9,13,17... So use the fórmula of the n-term ti find how MANY seats there are in row 20

OpenStudy (isaidavila):

Got it @zasharra

OpenStudy (zasharra):

is it 276?

OpenStudy (isaidavila):

I can't remember the formula. Can you give me the fórmula?

OpenStudy (zasharra):

20*10=200 4*19=76

OpenStudy (zasharra):

idk

OpenStudy (zasharra):

@Directrix

OpenStudy (zasharra):

276?

OpenStudy (isaidavila):

I think is that Tn=nr+to

OpenStudy (zasharra):

how will the equation go?

OpenStudy (isaidavila):

T(20)=20(4)-3

OpenStudy (isaidavila):

And the equation can be Tn=4n-3 for every row

OpenStudy (zasharra):

so am i right abut 276 seats on row 20/

OpenStudy (math&ing001):

number of seats depends on the number of the row, here's how it goes first row r(1)=10 r(2)= 10+4 r(n)= 10 +4(n-1)

OpenStudy (math&ing001):

How did you get 276 ?

OpenStudy (zasharra):

20*10=200 4*19=76

OpenStudy (zasharra):

then woulnt it be 10 14 18 ...so forth?

OpenStudy (zasharra):

guys plz help idk this

OpenStudy (isaidavila):

But how do you knlw row 1 has 10 seats. It says that CAN HAVE 10 seats

OpenStudy (math&ing001):

Exactly so for 20 it would be 10 +4(20-1) = 86

OpenStudy (zasharra):

it's in the question

OpenStudy (isaidavila):

But can have not have

OpenStudy (zasharra):

and @math&ing001 is that the final answer/

OpenStudy (math&ing001):

Yes and this is the equation to predict the seats in any row r(n)= 10 +4(n-1)

OpenStudy (zasharra):

20*4= 80 how u get 86?

OpenStudy (math&ing001):

why 20*4 they say "4 more seats than the row before" so it addition not multiplication it's 10+4*19

OpenStudy (zasharra):

yea you're right.. THANK YOU SO MUCH. I medal-ed u by the way!!

OpenStudy (math&ing001):

You're VERY WELCOME =3